OpenStudy (anonymous):

HELP PLEASE DOUBLE CHECK MY ANSWER!!! The quadratic functions f(x) and g(x) are described as follows: f(x) = −8x2 + 7 x g(x) 0 0 1 2 2 6 3 2 4 0 Which of the following statements best compares the maximum value of the 2 functions? It is the same for both functions. f(x) has a greater maximum value than g(x). g(x) has a greater maximum value than f(x). The maximum values cannot be determined. I chose A is that correct?

Vocaloid (vocaloid):

no, A is not correct let's just look at f(x). what is the maximum value?

Vocaloid (vocaloid):

not quite, let's take a look at f(x) f(x) = −8x2 + 7 what is the biggest possible value of f(x)?

OpenStudy (anonymous):

would it be 8?

Vocaloid (vocaloid):

not quite, the biggest possible value of f(x) occurs when x = 0 what is the value of f(x) when x = 0?

Vocaloid (vocaloid):

f(x) = -8x^2 + 7 what is f(x) when x = 0?

Vocaloid (vocaloid):

substitute 0 for x in the equation... f(x) = -8x^2 + 7 = -8(0)^2 + 7 = ?

OpenStudy (anonymous):

7

Vocaloid (vocaloid):

right, so the maximum value of f(x) is 7 now look at g(x). what is the maximum value of g(x)?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

so B ?

Vocaloid (vocaloid):

yes

OpenStudy (anonymous):

thank you